Suppose a 5-card hand is dealt to each of 2 players on a team. What is the expected total number of aces the team has?
Answer #1

total number of cards =52

total number of cards team has =10

number of aces in a 52 card deck =4

therefore expected total number of aces the team has =4*10/52 =40/52 =10/13